CRY+VIS Interface Meeting
Date: 2021/6/25 13:00-13:55
Participants: Takahashi, Hirata, Miyo, Kimura, Yokozawa, Ikeda, Node, Ushiba, Tomaru, Aso
- IX: MN was recovered. TF will be measured precisely in the next week.
- IY: Rough alignment of TM was tried. The ballast weight on the PF is too small to take a balance. Additional spring will be implemented on the BF on 7/28. Adjustment of ballast on the PF was scheduled on 7/29.
- EX: Alignment for the green beam is planed in the week of 7/5. OL damp in PR3, Yaw damp, and Oplev are necessary.
- EY: Mirror cleaning in Mitaka is on going. Though 24h x2 cleaning was applied, the stain remains. Still continue through the week end. Splay type FC will be tried.
- Corrugate tube fixing jig: Delivery and implement are scheduled on 7/12-13.
- TM height setting: Next target is ITMY.
- Commissioning in air: X-arm.
- Commissioning in vacuum:
- Overall commissioning:
- Acceptance check: Group work is set every Friday.
- PF actuator: After SC.
- Jig for FC painting:
- Separation of rack in 2F: On the maintenance days in July.
- Disassembly of PRM/PR2 assembly frame: This is scheduled on 6/28-29.
- New vacuum pumps for V-chambers: Check the port location and interference of cabling.
- Extension of cables for picomotor drivers: Test was planed in EYV.
Participants: Takahashi, Hirata, Miyo, Sato, Fabian, Sugiyama, Aso, Terrence, Tanaka, Ishizaki
This is generated by python script in /users/Miyo/dropbox/Script/scrapping/main.py.
- Assembled six accelerometers for IY and EX. The original capacitance sensors were replaced to the LVTDs.
- Designing the adapter plate for H4 geophones and the protection cover for accelerometers.
I wrote the Type-B section of the O3GK paper.
- With Takahashi-san we moved the ballast masses on the IP to have a more uniform distribution with the hope of reducing the asymmetry in the IP-L and IP-T resonance frequencies. It did not work. On the contrary, the difference increased: it went from 24 to about 30 mHz.
- Then we added 4.7 kg and the frequencies went down to 63 and 92 mHz.
- The temperature was 23.8 degrees, which is high. The frequencies will decrease when it becomes 23 degrees again.
- L and T look strongly coupled. Maybe similar to what Terrence observed in SRM.
We measured the IP transfer functions.
- The resonance frequencies in L and T are both close to 70 mHz.
- They cannot be distinguished from each other but they below 80 mHz.
- From measurements made this morning, the three legs seem equally stiff.
- The IP resonance frequencies in L and T had an asymmetry of about 23 mHz and one of them was very low: 70 and 47 mHz in L and T respectively.
- We removed 1.1 kg and redristributed 3.8 kg: the frequencies became 78 and 55 mHz.
- Later we removed 1.2 kg and they changed to 83 and 63 mHz.
- We were not able to remove or decrease the asymmetry.
- From measurements made this morning, it seems that the IP-H1 leg is stiffer than the others.
I have modeled the LVDT and geophone sensor noises. These will be useful for designing sensor correction filter, blending filters, and damping filters. Results here.
- I have also inter-calibrated the geophones with the LVDTs. Not posted yet. Will post after getting the geophone sensing matrix.
- I attempted to align the Geophones to the LVDTs but I have difficulties:
- I aligned the actuators to the LVDTs at 1 Hz, and tried to align the geophones to the actuators at 1 Hz. This way we can align the LVDTs and the geophones.
- But the geophones have a strange phase response at 1 Hz. This is probably due to poor calibration filter, but I am inclined not to fix this now.
- I am trying a different frequency, i.e. 0.2 Hz.
- The optical lever will become important soon. But the signals were zeroed out. Is it installed?
- Install the supporting beams of the BS outer frame.
- In BS, SR2 and SR3 lock the IP and F0 to measure noise.
- Measure parameters for the acceptance check.
- The green laser reflected by SR3 reached the X End.
- No major ajustments to the payload were necessary.
- It only required adjustment in yaw, which was done with the BF Damper first and later offloaded to the traverser.
- This likely means we left the payload in good condition.
- Type-A team: Adjustement of IYV
- Type-B team: Remaining work
IMC suspensions are ready in air  will finish 2021/04/21.
PRs are ready in air  will finish 2021/07/09.
ETMX is ready in air  will finish 2021/08/25.
ITMX is ready in air  will finish 2021/08/31.
ETMY is ready in air  will finish 2021/09/16.
ITMY is ready in air  will finish 2021/11/08.
BS is ready in air  will finish 2021/11/24.
SRs are ready in air  will finish 2021/12/20.
Output suspensions are ready in air  will finish 2021/12/21.
TMSY is ready in air  will finish 2021/??/??.
TMSX is ready in air  will finish 2021/??/??.
- Real-time model modification [Terrence]
- Two requests (Of course LIGO doesn't have them)
- Sensor correction path for the optical levers, this allows us to estimate the control performance without the interferometer. The seismometer signal exists in the model, I see no reason why we don't include that path.
- There's LVDTALIGN for aligning the LVDTs, which is good. But then I just realized the screen for ACCALIGN exists but they are all white. Can we at least add this for aligning the geophones/accelerometers?
- Two requests (Of course LIGO doesn't have them)
The week from 6/28
- Takahashi: 6/28-7/2
- Sato: 6/28-7/2
- Hirata: 6/28-7/2
- Yano: 6/28-7/2
The week from 7/5
- Takahashi: 7/5-9 (KIW8 7/7-9)
- Sato: 7/5-9
- Hirata: 7/5-9
- Yano: 7/5-9